OpenStudy (anonymous):

Find the least common multiple. X^2+8x, X^2-8x

OpenStudy (zzr0ck3r):

you need to factor both of the things \[(x^2+8x)=x(x+8)\\(x^2-8x)=\large?\]

OpenStudy (zzr0ck3r):

@grimmjow no looking around when im trying to help:)